# Dyonic Taub-NUT-AdS Black Branes: Thermodynamics and Phase Diagrams

### v1

@@ 3.2.1 Case I: $\alpha=0$
For $\alpha=0$, the $n-T$ phase diagram will only depend on the value of $\widetilde{Q}_{m}$. Generally, we have two cases;
@@
Figure 3.1: The temperature of the horizon of the black brane as a function of its horizon radius in case I.I for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=0$). The graph shows that for $n\leq\frac{\sqrt{3}}{3}l\phi_{e}$ the is only one possible black brane solution at any temperature, while for $n>\frac{\sqrt{3}}{3}l\phi_{e}$ there are two possible black brane solutions for $T_{h}\geq T_{min}$ and no possible solutions for $T_{h}<T_{min}.$
As shown in Fig.(3.1), if |$\absolutevalue{n}\leq\frac{\sqrt{3}}{3}l\phi_{e}$, the temperature is a monotonic function of $r_{h}$, meaning that only one black brane can exist at any temperature. Also, there is always an extremal solution with a horizon radius $r_{e}$ given by
$r_{e}=\frac{\sqrt{12\left(l^{2}\,\phi_{e}^{2}-3n^{2}\right)}}{6}.$ (53)
However, if |$\absolutevalue{n}>\frac{\sqrt{3}}{3}l\phi_{e}$, we see that the temperature has a minimum, $T_{min}$, at $r_{h}=r_{min}$ where
$r_{min}=\frac{\sqrt{9n^{2}-3l^{2}\phi_{e}^{2}}}{3},$ (54)
@@
It follows that at $T>T_{min}$ there are two possible black branes, a big black brane, and a small black brane. There is only one possible black brane at $T=T_{min}$. Finally, there are no possible black branes at $T<T_{min}$.
The Schwarzschild AdS solution [1] and some cases of the grand canonical ensemble of Reissner-Nordström AdS solution [4, 3] also have a minimum value for $T_{h}$. However, these cases always have a pure thermal solution that exists at any temperature which shares the same asymptotic metric as the black hole solutions. Our case lacks such a pure thermal solution. In fact, there is no known solution with the same asymptotic metric as the Taub-NUT dyonic AdS black brane metric. This means that in our case there are no known solutions that can exist at $T<T_{min}$. An educated guess would be that for $T<T_{min}$ the system will transit to a soliton-like solution, along the line that was presented in [42]. However, further investigation is required to confirm this assumption.
@@
  • For |$\absolutevalue{n}>\frac{\sqrt{3}}{3}l\phi_{e}$ and $T<T_{min}$, there are no possible solutions.
  • For |$\absolutevalue{n}>\frac{\sqrt{3}}{3}l\phi_{e}$ and $T=T_{min}$, the only possible solution is one black brane solution, and it is thermally stable.
  • For |$\absolutevalue{n}>\frac{\sqrt{3}}{3}l\phi_{e}$ and $T>T_{min}$, there are two possible black brane solutions, the small black brane is thermally unstable while the big black brane is thermally stable.
  • For |$\absolutevalue{n}\leq\frac{\sqrt{3}}{3}l\phi_{e}$, the only possible solution is one black brane solution, and it is thermally stable.
@@
To further investigate the different cases and the phases they contain, the grand potential density of the solution as a function of its horizon temperature is plotted for different values of $n$. This is shown n Fig. 3.2
(a) |$\absolutevalue{n}<\frac{\sqrt{3}}{3}l\phi_{e}$
(b) |$\absolutevalue{n}=\frac{\sqrt{3}}{3}l\phi_{e}$
(c) |$\absolutevalue{n}>\frac{\sqrt{3}}{3}l\phi_{e}$
Figure 3.2: The grand potential density of the black brane as a function of its horizon temperature in case I.I for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=0$). The graphs show that for $n\leq\frac{\sqrt{3}}{3}l\phi_{e}$ the grand potential of the black brane is always lower than that of the extremal black brane, while for $n>\frac{\sqrt{3}}{3}l\phi_{e}$ there is no extremal solution, and the stable black brane always have a lower grand potential than the unstable black brane.
From Fig. 3.2 we can see that for |$\absolutevalue{n}<\frac{\sqrt{3}}{3}l\phi_{e}$ the non-extremal black brane always has a lower potential density than the extremal one, making it the thermodynamically preferable solution. For |$\absolutevalue{n}>\frac{\sqrt{3}}{3}l\phi_{e}$ we can see the absence of an extremal solution. We can also see that no black brane can exist for $T<T_{min}$. On the other hand, for $T>T_{min}$ two black branes exist, with the thermally stable one always having a lower potential density than the non-stable one.
We are now fully acquainted with the required information to draw the $n-T$ phase diagram for this case. This is shown in Fig. 3.3. We can see from the figure that there is only one stable black brane everywhere except for the gray-shaded area where no black brane can exist.
Figure 3.3: The $n-T$ phase diagram for case I.I. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=0$). The graph shows that there exists only one phase everywhere except for the shaded gray area which marks where $T_{h}<T_{min}(n)$ and hence no black hole can exist in it.
@@
Due to the non-vanishing $\widetilde{Q}_{m}$, an additional $-\frac{1}{r_{h}^{3}}$ term is present in the temperature formula. Fig. 3.4 shows that regardless the value of $n$, there always exists an extremal black brane with horizon radius $r_{e}$ given by
[…]
Figure 3.4: The temperature of the horizon of the black brane as a function of its horizon radius in case I.II for different values of $n$. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=60$).
However, we either get three branches or one branch of the black brane solution depending on the value of $n$. We can also see that there are two critical values of $n$
$n_{c_{1}}=-\frac{\sqrt{18l\widetilde{Q}_{m}+3l^{2}\phi_{e}^{2}}}{3},\quad n_{c_{2}}=\frac{\sqrt{18l\widetilde{Q}_{m}+3l^{2}\phi_{e}^{2}}}{3},$ (58)
@@
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T<T_{min}$ or $T>T_{max}$, the only possible solution is one black brane solution, and it is thermally stable.
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T\in(T_{min},T_{max})$, there are three possible black brane solutions, the small and the big black branes are thermally stable while the middle one is thermally unstable
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T=T_{min}$ or $T=T_{max}$, there are two possible black brane solutions, and both of them are thermally stable.
  • For $n\in[n_{c_{1}},n_{c_{2}}]$, the only possible solution is one black brane solution, and it is thermally stable.
@@
As before, we first plot the grand potential density of the solution as a function of its temperature, as it contains all the necessary information to explore the phase diagram of the solution. This is shown in Fig. 3.5.
@@
Figure 3.5: The grand potential density of the black brane as a function of its horizon temperature in case I.II for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=60$).
From Fig. 3.5, we can immediately notice the swallowtail behavior of the grand potential density for $n<n_{c_{1}}$ and $n>n_{c_{2}}$, signifying a first-order phase transition of the system. The grand potential densities of the small and the big black branes cross at $T=T_{tr}$, which marks a first-order phase transition between the two black branes. Also, we can notice that the unstable middle black brane is never thermodynamically preferred. Moreover, for any value of $n$ and non-zero temperatures, there is always a solution with grand potential density lower than the extremal black brane, meaning it is never thermodynamically preferred. Summarizing, we have
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T<T_{min}$ or $T>T_{max}$, the only stable solution is one black brane solution.
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T\in[T_{min},T_{tr})$, there are two stable black brane solutions, the small black brane is thermodynamically preferred while the big black brane is a metastable state.
@@
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T\in(T_{tr},T_{max}]$, there are two stable black brane solutions, the big black brane is thermodynamically preferred while the small black brane is a metastable state.
@@
  • For $n\in[n_{c_{1}},n_{c_{2}}]$, the only stable solution is one black brane solution.
Figure 3.6: The $n-T$ phase diagram for case I.II. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=60$).
From Fig. 3.6 and all the subsequent phase diagrams, we can see that no phase transitions or critical points exist along the line $n=0$ regardless of the values of the other thermodynamic quantities. This confirms that a non-zero NUT charge is the reason for these behaviors, as they disappear when $n$ vanishes. It is also important to stress that this is the first study to report first-order phase transitions and critical behavior when studying the regular thermodynamics of black holes with flat horizons without adding extra terms to the gravitational action.
3.2.2 Case II: $0<\absolutevalue{\alpha}\leq\frac{\sqrt{3}}{3}$
The non-vanishing $\alpha$ will make this case even more rich and understandably more complicated. With all of our thermodynamic quantities non-vanishing, the horizon temperature will take the general form in Eq.(51). Following the same approach, we divide this case into sub-cases depending on the value of $\widetilde{Q}_{m}$.
Case II.I: |$\absolutevalue{\widetilde{Q}_{m}}<\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$
@@ 3.2.2 Case II: $0<\absolutevalue{\alpha}\leq\frac{\sqrt{3}}{3}$
[…]
and depending on the value of $n$ we either have one branch of the solution or three branches of the solution with two critical values of $n$,
$n_{c_{1}}=-\frac{3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}}$
$n_{c_{2}}=\frac{3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}}$ (62)
@@
Figure 3.7: The temperature of the horizon of the black brane as a function of its horizon radius in case II.I for different values of $n$. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=30$).
However, as we can see from Fig. 3.7, unlike case I.II, $n_{c_{1}}\neq-n_{c_{2}}$, $r_{h(c_{1})}\neq r_{h(c_{2})}$, and $T_{c_{1}}\neq T_{c_{2}}$, except for the special case when $\widetilde{Q}_{m}=0$, making the $n-T$ phase diagram, Fig.(3.8), not symmetric in general under $n\longrightarrow-n$ except for $\widetilde{Q}_{m}=0$ as we mentioned.
Figure 3.8: The $n-T$ phase diagram for case II.I. Values used to draw this graph are ($l=1$, $\phi_{e}=15$,$\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=30$).
Varying |$\absolutevalue{\widetilde{Q}_{m}}$ | within the range $[0,\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}})$ does not change the characteristics of the phase diagram but affects its shape. Decreasing |$\absolutevalue{\widetilde{Q}_{m}}$ | shrinks the interval $[T_{c_{1}},T_{c_{2}}]$ until $T_{c_{1}}=T_{c_{2}}$ at $\widetilde{Q}_{m}=0$. On the other hand, increasing |$\absolutevalue{\widetilde{Q}_{m}}$ | widens the interval $[T_{c_{1}},T_{c_{2}}]$ with $T_{c_{2}}\longrightarrow 0$ as |$\absolutevalue{\widetilde{Q}_{m}}\longrightarrow\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$.
Case II.II: |$\absolutevalue{\widetilde{Q}_{m}}=\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$
@@
Figure 3.9: The temperature of the horizon of the black brane as a function of its horizon radius in case II.II for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=45$).
Comparing this case with the previous one, we find that $n_{c_{1}}$ is the only critical point with
$n_{c}=\alpha^{2}l^{2}\phi_{e}^{3}\left(\sqrt{3}+3\alpha\right)$ (66)
@@
$n_{c_{2}}$ is no longer a critical point. However, it still separates between a three-branched solution and a one-branched solution for $T_{h}(r_{h})$. We will call it $n_{l}$, where
$n_{l}=\alpha^{2}l^{2}\phi_{e}^{3}\left(\sqrt{3}-3\alpha\right)$ (68)
The reason $n_{l}$ is not a critical value of $n$ is that, as we can see from Fig. 3.9, at $n=n_{l}$, the $T_{h}-r_{h}$ curve doesn’t have a saddle point. Instead, the horizon temperature of the black brane becomes a linear function of its horizon radius, $T_{h}=\frac{3r_{h}}{4\pi}$. That results in one of the phases being disjointed from the rest of the phases as seen in the phase diagram in Fig. 3.10.
Figure 3.10: The $n-T$ phase diagram for case II.II. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=45$).
Case II.III: |$\absolutevalue{\widetilde{Q}_{m}}>\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$
@@
Figure 3.11: The temperature of the horizon of the black brane as a function of its horizon radius in case II.III for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=60$). Here we split the graph into two for clarity.
Similar to the previous case, we find two critical values for $n$,
$n_{c_{1}}=-\frac{3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}}$
$n_{c_{2}}=\frac{-3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}},$ (69)
@@
with $T_{h}(r_{h})$ has one black brane branch for $n\in[n_{c_{1}},n_{c_{2}}]$ and three branches for $n<n_{c_{1}}$ or $n>n_{c_{2}}$, except when $n=\frac{\widetilde{Q}_{m}}{\alpha\phi_{e}}$. We will give this value of $n$ a special name, $n_{hp}$, where hp stands for Hawking-Page. The reason for this is that for $n=n_{hp}$ the solution behaves exactly like a Schwarzschild anti-de Sitter solution described in [1]. At this value of $n$, the horizon temperature has a minimum value $T_{min}$;
@@
For $T>T_{min}$, one thermally stable and one thermally unstable black brane exist. For $T=T_{min}$, one thermally stable black brane exists, and no black branes exist for $T<T_{min}$. In fact, as we discussed in case I.I, no known solution can exist at this range of temperatures. To check the thermal preferability of the phases, we plot $\Omega(T_{h})$, Fig 3.13.
@@
Figure 3.13: The grand potential density of the black brane as a function of its horizon temperature in case II.III for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=60$).
@@
  • For $\,n<n_{c_{1}}$, $\;n\in(n_{c_{2}},n_{hp})$, or $\;n>n_{hp}\,$ and $\,T<T_{min}$ or $T>T_{max}\,$, the only stable solution is one black brane solution.
  • For $\,n<n_{c_{1}}$, $\;n\in(n_{c_{2}},n_{hp})$, or $\;n>n_{hp}\,$ and $T\in[T_{min},T_{tr})$, there are two stable black brane solutions. The small black brane is thermodynamically preferred, while the big black brane is a metastable state.
@@
  • For $\,n<n_{c_{1}}$, $\;n\in(n_{c_{2}},n_{hp})$, or $\;n>n_{hp}\,$ and $T\in(T_{tr},T_{max}]$, there are two stable black brane solutions. The big black brane is thermodynamically preferred while the small black brane is a metastable state.
@@
  • For $n\in[n_{c_{1}},n_{c_{2}}]$, the only stable solution is one black brane solution.
  • For $n=n_{hp}$ and $T<T_{min}$, there are no possible solutions.
  • For $n=n_{hp}$ and $T\geq T_{min}$, the only stable solution is a one black brane solution.
@@
Figure 3.14: The $n-T$ phase diagram for case II.III. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=60$). The gray line at $n=n_{hp}$ shows the region where no black brane can exist.
3.2.3 Case III: |$\absolutevalue{\alpha}>\frac{\sqrt{3}}{3}$

### v2

@@ 3.2.1 Case I: $\alpha=0$
For $\alpha=0$, the nTn-T phase diagram will only depend on the value of $\widetilde{Q}_{m}$. Generally, we have two cases;
@@
Figure 3.1: The black brane horizon temperature as a function of the horizon radius in case I.I for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=0$). The graph shows that for $n\leq\frac{\sqrt{3}}{3}l\phi_{e}$ the is only one possible black brane at any temperature, while for $n>\frac{\sqrt{3}}{3}l\phi_{e}$ there are two possible black brane phases for $T_{h}\geq T_{min}$ and no possible black brane phase for $T_{h}<T_{min}.$
[…]
$r_{e}=\frac{\sqrt{3\left(l^{2}\,\phi_{e}^{2}-3n^{2}\right)}}{3}.$ (62)
[…]
$r_{min}=\frac{\sqrt{3\left(3n^{2}-l^{2}\phi_{e}^{2}\right)}}{3},$ (63)
@@
It follows that at $T>T_{min}$ there are two possible black branes, a big black brane, and a small black brane, there is only one possible black brane at $T=T_{min}$, and finally, if $T<T_{min}$ no black brane can exist in thermal equilibrium, since the black brane horizon temperature cannot fall below $T_{min}$. Therefore, the black brane cannot be considered as a phase at this range of temperatures.
The Schwarzschild AdS solution [35] and some cases of the grand canonical ensemble of Reissner-Nordström AdS solution [17, 16] also have a minimum value for $T_{h}$. However, these cases always have a pure thermal solution that exists at any temperature which shares the same asymptotic metric as the black hole. Our case lacks such a pure thermal solution. In fact, there is no known solution with the same asymptotic metric and have the same conserved charges as the Taub-NUT dyonic AdS black brane solution. This means that in our case there are no known solution that can exist at $T<T_{min}$. An educated guess would be that for $T<T_{min}$ the system will transit to a soliton-like solution, along the line that was presented in [50]. However, further investigation is required to confirm this assumption.
@@
[…]
@@
To further investigate the different cases and the phases they contain, the grand potential density of the solution as a function of its horizon temperature is plotted for different values of $n$. This is shown in Fig. 3.2
[…]
Figure 3.2: The grand potential density of the black brane as a function of the horizon temperature in case I.I for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=0$). The graphs show that for $n\leq\frac{\sqrt{3}}{3}l\phi_{e}$ the grand potential of the black brane is always lower than that of the extremal black brane, while for $n>\frac{\sqrt{3}}{3}l\phi_{e}$ there is no extremal black brane, and the stable black brane always have a lower grand potential than the unstable black brane.
[…]
We are now fully acquainted with the required information to draw the nTn-T phase diagram for this case. This is shown in Fig. 3.3. We can see from the figure that there is only one stable black brane phase everywhere except for the gray-shaded area, in which $T<T_{min}(n)$, where black brane cannot exist as a phase.
Figure 3.3: The nTn-T phase diagram for case $I.$ The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=0$). The graph shows that there exists only one phase everywhere except for the shaded gray area which marks where $T<T_{min}(n)$ and hence no black brane can exist in thermal equilibrium.
@@
Due to the non-vanishing $\widetilde{Q}_{m}$, an additional $\frac{1}{r_{h}^{3}}$ term is present in the temperature formula. Fig. 3.4 shows that regardless the value of $n$, there exists an extremal black brane with horizon radius $r_{e}$ given by
$r_{e}=\frac{\sqrt{6\left(l^{2}\,\phi_{e}^{2}-3n^{2}\right)\,+\,6\,\sqrt{\left(l^{2}\,\phi_{e}^{2}-3n^{2}\right)^{2}\,+\,12l^{2}\widetilde{Q}_{m}^{2}}}}{6}.$ (66)
Figure 3.4: The black brane horizon temperature as a function of the horizon radius in case I.II for different values of $n$. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=60$).
However, we either get a monotonic or non-monotonic $T_{h}(r_{h})$ depending on the value of $n$. with two critical points
$n_{c_{1}}=-\frac{\sqrt{18l\widetilde{Q}_{m}+3l^{2}\phi_{e}^{2}}}{3},\qquad n_{c_{2}}=\frac{\sqrt{18l\widetilde{Q}_{m}+3l^{2}\phi_{e}^{2}}}{3},$ (67)
@@
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T<T_{min}$ or $T>T_{max}$, the only existing phase is one black brane, and it is thermally stable.
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T\in(T_{min},T_{max})$, there are three possible black brane phases, the small and the big black branes are thermally stable while the middle one is thermally unstable.
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T=T_{min}$ or $T=T_{max}$, there are two possible black brane phases, and both of them are thermally stable.
  • For $n\in[n_{c_{1}},n_{c_{2}}]$, the only existing phase is one black brane, and it is thermally stable.
@@
As before, we first plot the grand potential density of the solution as a function of the horizon temperature, as it contains all the necessary information to explore the phase diagram. This is shown in Fig. 3.5.
@@
Figure 3.5: The grand potential density of the black brane as a function of the horizon temperature in case I.II for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=60$).
From Fig. 3.5, we can immediately notice the swallowtail behavior of the grand potential density for $n<n_{c_{1}}$ and $n>n_{c_{2}}$, signifying a first-order phase transition of the system. The grand potential densities of the small and the big black branes cross at $T=T_{tr}$, which marks a first-order phase transition between the two black branes. Also, we can notice that the unstable middle black brane is never thermodynamically preferred. Moreover, for any value of $n$ and non-zero temperatures, there is always a phase with grand potential density lower than the extremal black brane, meaning that extremal black brane is also never thermodynamically preferred. Summarizing, we have
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T<T_{min}$ or $T>T_{max}$, the only existing phase is one black brane.
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T\in[T_{min},T_{tr})$, there are two stable black brane phases, the small black brane is thermodynamically preferred while the big black brane is metastable.
@@
  • For $n<n_{c_{1}}$ or $n>n_{c_{2}}$, and $T\in(T_{tr},T_{max}]$, there are two stable black brane phases, the big black brane is thermodynamically preferred while the small black brane is metastable.
@@
  • For $n\in[n_{c_{1}},n_{c_{2}}]$, the only existing phase is one black brane.
Figure 3.6: The nTn-T phase diagram for case I.II. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=0$, $\widetilde{Q}_{m}=60$).
From Fig. 3.6 and all the subsequent phase diagrams, we can see that no phase transitions or critical points exist along the line $n=0$ regardless of the values of the other thermodynamic quantities. This confirms that a non-zero NUT charge is the reason for these behaviors, as they disappear when $n$ vanishes. It is also important to stress that this is the first study to report rich phase structure with first-order phase transitions and critical behavior when studying the regular thermodynamics of black holes with flat horizons without adding extra terms to the gravitational action.
[…]
The non-vanishing $\alpha$ will make this case even more rich and understandably more complicated. With all of our thermodynamic quantities non-vanishing, the horizon temperature will take the general form in Eq.(51). Following the same approach and in light of our analysis of the critical point, we divide this case into sub-cases depending on the value of $\widetilde{Q}_{m}$.
[…]
@@
$r_{e}=\frac{\sqrt{6\left(l^{2}\,\phi_{e}^{2}-3n^{2}\right)\,+\,6\,\sqrt{\left(l^{2}\,\phi_{e}^{2}-3n^{2}\right)^{2}\,+\,12l^{2}\left(\alpha n\phi_{e}-\widetilde{Q}_{m}\right)^{2}}}}{6},$ (70)
and depending on the value of $n$ we either have monotonic or non-monotonic $T_{h}(r_{h})$ with two critical values of $n$,
$n_{c_{1}}=-\alpha l\phi_{e}-\frac{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3},$
$n_{c_{2}}=\alpha l\phi_{e}+\frac{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}{3},$ (71)
@@
Figure 3.7: The black brane horizon temperature as a function of the horizon radius in case II.I for different values of $n$. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=30$).
However, as we can see from Fig. 3.7, unlike case I.II, $n_{c_{1}}\neq-n_{c_{2}}$, $r_{h(c_{1})}\neq r_{h(c_{2})}$, and $T_{c_{1}}\neq T_{c_{2}}$, except for the special case when $\widetilde{Q}_{m}=0$, giving rise to a non-symmetric phase diagram under $n\longrightarrow-n$ in general, Fig. 3.8, except for $\widetilde{Q}_{m}=0$ as we mentioned.
Figure 3.8: The nTn-T phase diagram for case II.I. Values used to draw this graph are ($l=1$, $\phi_{e}=15$$\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=30$).
[…]
@@
Figure 3.9: The black brane horizon temperature as a function of the horizon radius in case II.II for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=45$).
Following our analysis for the critical points we find that we have only one critical point.
$n_{c}=-\frac{l\phi_{e}}{3}\left(6\alpha+\sqrt{3}\right),$ (75)
@@
$n_{c_{2}}$ is no longer a critical point. However, it still separates between a monotonic and a non-monotonic $T_{h}(r_{h})$. We will call it $n_{l}$, where
$n_{l}=\frac{l\phi_{e}}{\sqrt{3}}$ (77)
The reason $n_{l}$ is not a critical value of $n$ is that $r_{h(c_{2})}=0$ under the given conditions. This is indicated from Fig. 3.9 where at $n=n_{l}$, the $T_{h}-r_{h}$ curve doesn’t have any saddle point. Instead, the horizon temperature of the black brane becomes a linear function of its horizon radius, $T_{h}=\frac{3r_{h}}{4\pi l^{2}}$. That results in one of the phases being disjointed from the rest of the phases as seen in the phase diagram in Fig. 3.10.
Figure 3.10: The nTn-T phase diagram for case II.II. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=45$).
[…]
@@
Figure 3.11: The black brane horizon temperature as a function of the horizon radius in case II.III for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=60$). Here we split the graph into two for clarity.
Referring to our analysis, only two critical points will survive, $n_{c_{1}}$ and $n_{c_{2}}$,
$n_{c_{1}}=-\alpha l\phi_{e}-\frac{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3},$
$n_{c_{2}}=-\alpha l\phi_{e}+\frac{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3},$ (78)
@@
with $T_{h}(r_{h})$ being monotonic for $n\in[n_{c_{1}},n_{c_{2}}]$ and non-monotonic for $n<n_{c_{1}}$ or $n>n_{c_{2}}$, with three branches, except for $n=\frac{\widetilde{Q}_{m}}{\alpha\phi_{e}}$ where the solution behaves exactly like a Schwarzschild anti-de Sitter solution described in [35]. Therefore, We will give this value of $n$ a special name, $n_{hp}$, where hp stands for Hawking-Page. At this value of $n$, the horizon temperature has a minimum value $T_{min}$;
@@
For $T>T_{min}$, one thermally stable and one thermally unstable black brane exist. For $T=T_{min}$, only one thermally stable black brane phase exists, and no black branes phase exist for $T<T_{min}$. In fact, as we discussed in case I.I, no known solution with the same asymptotic metric can exist at this range of temperatures. To check the thermal preferability of the phases, we plot $\Omega(T_{h})$, Fig 3.13.
@@
Figure 3.13: The grand potential density of the black brane as a function of the horizon temperature in case II.III for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=60$).
@@
  • For $n<n_{c_{1}}$, $n\in(n_{c_{2}},n_{hp})$, or $n>n_{hp}\,$ and $T<T_{min}$ or $T>T_{max}\,$, the only stable phase is one black brane.
  • For $n<n_{c_{1}}$, $n\in(n_{c_{2}},n_{hp})$, or $n>n_{hp}\,$ and $T\in[T_{min},T_{tr})$, there are two stable black brane phases. The small black brane is thermodynamically preferred, while the big black brane is metastable.
@@
  • For $n<n_{c_{1}}$, $n\in(n_{c_{2}},n_{hp})$, or $n>n_{hp}\,$ and $T\in(T_{tr},T_{max}]$, there are two stable black brane phases. The big black brane is thermodynamically preferred while the small black brane is metastable.
@@
  • For $n\in[n_{c_{1}},n_{c_{2}}]$, the only stable phase is one black brane.
  • For $n=n_{hp}$ and $T<T_{min}$, there are no possible black brane phases.
  • For $n=n_{hp}$ and $T\geq T_{min}$, the only stable phase is a one black brane.
@@
Figure 3.14: The nTn-T phase diagram for case II.III. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=\frac{\sqrt{3}}{5}$, $\widetilde{Q}_{m}=60$). The gray line at $n=n_{hp}$ shows the region where black brane cannot exist as a phase.
[…]